OpenStudy (anonymous):

I ran into the following exercise: Define \(f:(0,c)\to\mathbb{R}\), for some \(c\in\mathbb{R}\) with \(00\). Moreover, let \(\epsilon=2c\delta\). Then, because \(0

OpenStudy (anonymous):

Does the fact that we are reducing the domain makes the function uniformly continuous? I don't think that f(x) = x^2 is UC in general. But I don't remember. Anyway, I can't see an error on your proof. Maybe someone with a sharper eye can see it.

OpenStudy (anonymous):

You are correct, \(f\) is not continuous over an unbounded domain. That's exactly the motivation behind this question.

OpenStudy (anonymous):

I forgot to mention the page no. It is page 2 ,example 6

OpenStudy (anonymous):

lol, sorry for my lack of knowledge but how do you read the definition of UC function?

OpenStudy (anonymous):

Oh, nice. I didn't know this was a canonical exercise. With that clarified, could I assert that \(f\) is uniformly continuous as \(c\to\infty\)? This would yield a contradiction, however.

OpenStudy (anonymous):

See eg 10 for your answer in the same pdf :D

OpenStudy (anonymous):

Gee, I need to learn how to read! Thanks, shivam. :)

OpenStudy (anonymous):

@pre-algebra , LOL, trust me, I have never encountered this stuff ever before. But still, I can understand something of what they wrote :D

OpenStudy (anonymous):

Suppose it is. Let \(\epsilon = 1\) Pick \(\delta > 0 \)(by uniform continuity) such that \( |x - y| < \delta\) implies \( |x^2 - y^2 | < 1 \) But \( |x^2 - y^2| = |x - y| * |x + y| \) and if we let x = n, y = n + \( \delta /2 \), for some \(n \in \mathbb{N}\), that implies that |x - y | = \( \delta/2 < \delta\) but \( |x^2 - y^2 | = d/2 * (2n + d/2)\) which is > 1 for a big enough n. So f cannot be uniformly continuous. I think this is a sufficient proof for \( f\) On a note, see that d is fixed, so the latter tends to infinity as n gets bigger.

OpenStudy (anonymous):

Typo, those d/2 should've been \(\delta/2 * (2n + \delta/2) \)

OpenStudy (anonymous):

And latter here refers to |x^2 - y^2|. Damn, my English is so ambiguous